Find the section of the conicoid (x/2) − (y 2/ 3 )= 2z by the plane x− 2y+ z = 1. What conic does this section represent? Justify your answer.


1
Expert's answer
2021-06-13T16:06:46-0400

Let us find the section of the conicoid "\frac{x^2}{2} −\frac{y^2}{3}= 2z"  by the plane x− 2y+ z = 1.


It follows that "\frac{x^2}{2} −\frac{y^2}{3}= 2(1-x+2y)," and therefore "\frac{x^2}{2} −\frac{y^2}{3}+2x-4y-2=0." The last equality equivalent to "3x^2-2y^2+12x-24y-12=0," and thus "3(x^2+4x)-2(y^2+12y)-12=0." It follows that "3(x+2)^2-2(y+6)^2=-48," and consequently the equation of the section is "\frac{(y+6)^2}{24}-\frac{(x+2)^2}{16}=1."

It follows that this section represents a hyperbola.